10 kl 72 k.ft 1061 8* 72 k.lt 8 kips 2 k/ft

Draw the shear and moment diagrams
Answer #1

tok 72 kift v 76 Let VB, VE be the vertical reactions at supports B, E respectively. Calculation of support reactions * VatVEshear force calculations (sf) ; (S.F) & 2 +8k +8-VE = +8 - 18.86 - - 10.86k cuj (SF)- (SF) = -10,86 k (iv) (S.Fc (just to the(iv) (B.m) (just to the better off) - 17.16 – 72 – – Kift. Vem)c = -8x(6+6+4) + Ve x[6+4) -72 – – 11.4 kift (vi) (B.MA=0 (ii)